Please help with this word problem:
A woman made a deposit of $183. If her deposit consisted of 83 bills, some of them one-dollar bills and the rest being five-dollar bills, how many one-dollar bills did she deposit?
thanks

Answer by Earlsdon(6103)

Let x = the number of $1.00 bills.
Then: x($1.00) + (83 - x)($5.00) = $183.00 Simplify and solve for x.
x + 415 - 5x = 183.00
415 - 4x = 183 Add 4x to both sides.
415 = 183 + 4x Subtract 183 from both sides.
232 = 4x Divide both sides by 4.
x = 58
The woman deposited 58 $1.00 bills
Check:
58($1.00) + (83 - 58)($5.00) = $58.00 + 25($5.00) = $58.00 + $125.00 = $183.00